OpenStudy (anonymous):

A line with slope -4/5 passes through the point (10, -7). What is y when x = -5?

sam (.sam.):

find slope equation

OpenStudy (anonymous):

I know I need to go over then down right from the point given

sam (.sam.):

y-y1=m(x-x1) y+7=-4/5(x-10) y=-4x/5 +8-7 y=-4x/5 +1

sam (.sam.):

when x=-5, y=-4x/5 +1 y=-4(-5)/5 +1 y=5

OpenStudy (anonymous):

ok i see